An investor invested a total of ​$800 in two mutual funds. One fund earned a 7​% profit while the other earned a 2% profit. If the​ investor's total profit was ​$26​, how much was invested in each mutual​ fund?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.
An investor invested a total of ​$800 in two mutual funds.
One fund earned a 7​% profit while the other earned a 2% profit.
If the​ investor's total profit was ​$26​, how much was invested in each mutual​ fund?
~~~~~~~~~~~~~~~~~

Let x be the amount in dollars invested at 7%;
then the amount invested at 2% is the reat 800-x dollars.


Write the total annual interest equation

    0.07x + 0.02*(800-x) = 26  dollars.


Simplify it and find x

    0.07x + 16 - 0.02x = 26

    0.07x - 0.02x = 26 - 16

         0.05x    =    16

             x    =    16/0.05 = 1600/5 = 320.


ANSWER.  $320 invested at 7%, the rest, or $800 - $320 = $480 invested at 2%.


CHECK.  0.07*320 + 0.02*480 = 32 dollars, the total annual interest.   ! correct !

Solved.



Answer by greenestamps(13200)



Here is an quick and easy informal method for solving any 2-part "mixture" problem like this.

(1) All $800 invested at 2% would earn $16 interest; all at 7% would earn $56 interest.
(2) The actual interest, $26, was 10/40 = 1/4 of the way from $16 to $56. ($16 to $56 is a difference of 40; $16 to $26 is a difference of 10; 10/40 = 1/4.)
(3) That means 1/4 of the total was invested at the higher rate.

1/4 of $800 is $200; so

ANSWERS: $200 was invested at 7% and the other $600 at 2%.

CHECK: .07(200)+.02(600) = 14+12 = 26